RAD Co.’s most recent free cash flow was $1,000,000. The firm is expected to grow free cash flow at a constant rate of 3% per year forever. RAD Co. has $2,000,000 of debt, $800,000 of preferred stock, and 100,000 shares outstanding. The Weighted Average Cost of Capital is 7%. Compute firm value and today’s stock price.
Answer #1

Current Free Cash Flow (FCF0) = $ 1000000, Perpetual Constatnt Growth Rate = 3 % and WACC = 7 %

FCF1 = 1000000 x 1.03 = $ 1030000

Firm Value = FCF1 / (WACC - Growth Rate) = 1030000 / (0.07 - 0.03) = $ 25750000

Less: Debt Value = $ 2000000

Less: Preferred Stock = $ 800000

Equity Value = $ 22950000

Number of Shares Outstanding = 100000

Price per Share = Stock Price = 22950000 / 100000 = $ 229.5